OpenStudy (anonymous):

5, 8, 11, 14, 17 .Which function can be used to find the xth number in this sequence?

OpenStudy (bahrom7893):

5+3=8 8+3=11 11+3=14

OpenStudy (bahrom7893):

so what function are you looking at?

OpenStudy (anonymous):

2x+3

OpenStudy (bahrom7893):

uhm sorry nope

OpenStudy (bahrom7893):

Or it would be: 2*5+3 = 13 2*13+3 = 29..

OpenStudy (jack1):

3x +5 where x= n-1...?

OpenStudy (jack1):

sweeeeeet ;)

OpenStudy (bahrom7893):

Ok I was taking the wrong approach. Jack's right

OpenStudy (anonymous):

The difference between two consecutive terms is always 3, so we can write : \[u_n=3n+a\] To find a we have \[u_1=a+3=5\] So : \[a=2\] Therefore : \[u_n=3n+2\]

OpenStudy (anonymous):

thanks